# InstructAttribute: Fine-grained Object Attributes editing with Instruction

### v1

@@ 4.2 Data Collection With MLLMs
Instructional training dataset generation. At this stage, we randomly select 200 pairs of source and target images where the object in the target images has varying attributes. Using GPT-4’s in-context learning, we prompt the model to identify the differences between each pair and how we can convert one image into another image. Based on its responses, we develop two types of editing instruction templates to enhance attribute editing diversity: (a) instructions for transforming the source image into the target image, and (b) instructions for generating target images with different attributes. The first type of instruction guides the model in altering intrinsic object attributes to entirely different ones, while the second enables transitions between different attributes of the same object. For color modification, we add a new category of editing instruction templates focused on generating target images within the same color hue. This instruction enhances the model’s ability to fine-tune colors within the same hue by adjusting brightness and saturation levels. For each pair of source-target images, we randomly select an instruction template to create the triple training dataset. An example of a dataset sample is shown in Fig. 7. More detailed information on the instruction templates is given in Appendix (c).
4.3 Instructional Supervised Training
@@
With the object attribute dataset, we adopt the instructional fine-tuning approach of InstructPix2Pix [2] to train our method to edit object attributes. This approach utilizes a pre-trained Latent Diffusion Model (LDM) [35] as its primary framework, implementing a two-stage synthesis process that performs diffusion and denoising in a compressed latent space while ensuring high-quality results. During training, a pre-trained autoencoder with an encoder $E$ and a decoder $D$ is used to first map the edited image $x$ to its latent representation $z=E(x)$. The model then learns a network $\theta$ to predict the noise added to the noisy latent $z_{t}$, conditioned on both the image $c_{I}$ and the text instruction $c_{T}$, where $c_{T}$ is concatenated with $z_{t}$ and $\epsilon(c_{I})$. The network $\theta$ , which is initialized with Stable Diffusion weights, is fine-tuned to minimize the latent diffusion objective:
@@
5 Attribute Editing Dataset
To construct the ATTRIBUTE dataset, initially, we collected class names from established datasets COCO [24], ImageNet [6], Open Images Dataset [22] and VisualGenome [21], resulting in a comprehensive list of more than 3,300 subjects, including foods, animals, plants, toys, clothing, transportations, daily necessities, natural objects , and more. Subsequently, for each subject, we generate the source images using 19 distinct prompt templates combined with randomly selected seeds. As altering a subject’s material inherently modifies its structure and texture, whereas color modification should minimally affect these properties, we divide the Attribute dataset into a color-editing dataset and a material-editing dataset. Once the datasets are generated, they are integrated and utilized to train the InstructAttribute model.
Color editing dataset. In this study, we focus on 43 colors that are frequently observed in daily life, systematically chosen from the families of red, orange, yellow, green, blue, purple, brown and neutral colors, and a detailed list of color names is given in the Appendix (d). In practice, we find that initially the cross-attention value amplifying ratio of the color descriptor to 5 and gradually reducing it by 0.1 per denoising step until it reaches 1.0 yields the optimal target color image success rate. Finally, we generate more than 638 K pairs of source-target color image pairs after the image selection process, and the distribution of different color images is given in Appendix (d). Through the instructional training dataset generation process, the color editing scenario, a triple dataset consisting of 1.5M input-output image pairs is generated, along with corresponding edit instructions.
Material editing dataset. we examine 14 materials commonly observed in daily life, including cotton, glass, marble, plastic, velvet, denim, lace, mesh, wood, fur, leather, metal, suede, and wool. For the target material subject generated, we find that initializing the cross-attention Value amplifying ratio at 10 of the material descriptor and progressively reducing it by 0.2 per denoising step until reaching 1.0 achieves the highest success rate for generating target material images. In total, we generate over 198 K source-target material image pairs after the image selection process, and the distribution of different material images is given in the Appendix (d). Using those image pairs, we create a triple dataset for the material editing scenario, consisting of 396K input-output image pairs and their corresponding editing instructions.
@@ 6 Experiment
Figure 8: Comparative results of attributes alteration with different instruction-based editing methods. As demonstrated, our method achieves the optimal balance when modifying the attributes of an object and preserving the its original structural integrity and background information
Figure 9: More Results of attribute modification. As shown, our method enables more precise color adjustment, performing color change within the same color hue, and facilitates a consistent material modification effect across various subjects.
@@
Baseline models. We make comparisons with the state-of-the-art text-guided image editing methods, including InstructPix2Pix (IP2P) [2], Instructdiffusion (InsDiff) [10], HIVE [14], MGIE [9], and MagicBrush [42]. To maintain reproducibility and fairness, we employed the default hyperparameters provided in the official implementations.
Qualitative evaluation. In the color alteration task, as demonstrated in Fig. 8 (a) and (c), our model produces the best color alteration result, while the methods IP2P and MGIE may alter background information, InsDiff, HIVE, and MagicBrush do not change the object color correctly and may change the structure of the object. For the material-changing task, as demonstrated in Fig. 8 (b) , methods such as IP2P, InsDiff, HIVE, MGIE and MagicBrush face challenges to perform the modification and make structural deformation, while our method achieves the optimal balance between material alteration and preservation of the original structure of the object. Furthermore, Fig. 9 (a) shows that our model is capable of accurate and precise color adjustments while maintaining the structural integrity of the subject. Fig. 9 (b) reveals that our method is capable of modifying the color of an object within the same hue by changing the saturation and brightness of the color. Fig. 9 (c) illustrates that our method ensures a consistent effect of material transformation on various subjects while preserving the structure of the object and the background information. More examples of color and material modification are given in Appendix (e).
@@
IP2P 16.8% 7.6% 8.8% 9.8%
InsDiff 4.0% 4.0% 4.8% 1.8%
HIVE 9.2% 8.6% 6.8% 7.0%
MGIE 6.2% 2.4% 3.8% 4.0%
MagicBrush 11.6% 4.6% 9.0% 6.8%
Ours 52.2% 72.8% 66.8% 70.6%
@@
Quantitative evaluation. Given the lack of publicly available datasets for evaluating material and color modification tasks, we developed two specialized datasets designed for quantitative analysis, addressing material and color modification tasks independently. For the material modification task, we randomly selected 100 subjects and generated 239 source images with StableDiffusion [35] checkpoints V1.4 with a random seed. For each source image, we modified the material to all 14 fine-grained materials, resulting in a total of 3,346 pairs of source and target images. For evaluation metrics, we used the DINO score [4] on a gray scale to measure semantic similarity, SSIM [41] to measure structural similarity, and used the CLIP Score (CS) [33] to measure similarity between the target image and its corresponding text description. For the color alteration task, since any object can undergo a color change, we enlarge the subject type to 659 and generate a source image for each subject using the StableDiffusion [35] checkpoints V1.4. For each source image, we used the SAM method [20] to generate the object mask and manually selected the best, resulting in 659 source images with the corresponding subject mask. For each source image, we modified the color to all 43 fine-grained colors, resulting in a total of 28,337 pairs of source and target images. For evaluation metrics, except for using the DINO score, SSIM, and CLIP score, we used LPIPS [43] in the background area to measure the preservation of background information. L1 loss of HVS of the object area between the target image and the corresponding pure color image to measure how well the color of the object has been changed. As can be seen in Table 2, for the color alteration task, our method performs best in preserving structure, background information, and the best at changing the object’s color compared to other instruct-based methods. For the material modification task, our method performs best in preserving the structure while having the highest text-image similarity after we changed the material of the subject.
@@
Human Preference Study. We conduct a human preference study to measure the performance of different methods asking: (1) which method produces the best attributes modification; (2) which method produces the best preserves the structure of the subject; (3) which method produces the best preserves the background; (4) which method produces the best overall attribute editing quality. We randomly sampled 100 source-target examples and let 20 annotators assess those images. The results of the human preference study are listed in Table 3. As can be seen, our method performs the best in all aspects, and the results are consistent with the quantitative experiments. For more detailed information, see the Appendix (f).
@@ 7 Conclusion
In the work, we propose an instruction-based model for object-level color and material modification, designed to allow precise and controlled changes to the visual attributes of objects while maintaining the integrity of their structure and surrounding context. First, we propose a training-free structure preservation and attribute amplification method, which generates diverse color and material variations for the same object without requiring additional training or fine-tuning. Next, leveraging our SPAA method, we generate the Attribute Dataset, the first large-scale, diverse dataset specifically designed for object color and material editing. This dataset contains a wide variety of objects, colors, and materials, providing a solid foundation for the domain of attribute modification. Finally, we train our instruction-based model on the Attribute dataset and achieve an optimal balance between precise attribute modification and preservation of the original object structure and background information. Limitations and Future Work. While our SPAA method produces plausible color and material variations of an object, we acknowledge that attribute modification may fail at times due to the mis-distributed cross-attention map. Currently, we address this issue by using LLaVA and Dino to filter out unsuitable target images, but this approach can sometimes exclude valid variations. In the future, our aim is to explore more sophisticated solutions to the cross-attention leakage problem, potentially through enhanced attention mechanisms or architectural innovations that explicitly constrain cross-attention to focus on the target object while minimizing interference from irrelevant regions.
[…]
Supplementary Material
8 Rationale
a) Pseudocode of SPAA algorithm
Algorithm 1 Object-level Color and material Editing Algorithm for a generated image
Input: A source prompt $P^{S}$, A target prompt $P^{T}$, an interesting attribute $A$ in $P_{T}$ Output: A edited image $I^{T}$
2:$z^{T}_{T}\leftarrow z^{T}_{S}$
3:for $t=T,T-1,...,1$ do
4:       $z^{T}_{t},M^{S}_{cross_{t}},M^{S}_{self_{t}}=\text{DM}(z^{S}_{t},P_{S},t)$
5:       $M^{T}_{cross_{t}}=\text{DM}(z^{T}_{t},P^{T},t)$
6:       $\hat{M}^{T}_{cross_{t}}=\text{Edit}(M^{S}_{cross},M^{T}_{cross_{t}},t)$
7:       $z^{T}_{t-1}=\text{DM}(z^{T}_{t},P^{T},t){M^{T}_{self}\leftarrow M^{S}_{self},M^{T}_{cross_{t}}\leftarrow\hat{M}^{T}_{cross_{t}}}$
8:end for
9:$I^{T}\leftarrow Generate(z^{T}_{0})$
10:return $I^{T}$
11:
12:if $A(j)=None$ then
13:       $\text{Edit}(M^{S}_{cross},M^{T}_{cross_{t}},t)_{i,j}=M^{T}_{cross_{i,j}}$
14:else
15:       $\text{Edit}(M^{S}_{cross},M^{T}_{cross_{t}},t)_{i,j}=M^{T}_{cross_{i,A_{j}}}$
16:end if
17:
18:if Token(j) is Attribute Discriptor then
19:       $\text{DM}(z^{T}_{t},P^{T},t)_{j}=M^{T}_{cross_{i,A_{j}}}*(ratio*V_{j})$
20:       $ratio=(1-\lambda)*ratio$
21:else
22:       $\text{DM}(z^{T}_{t},P^{T},t)_{j}=\text{DM}(z^{T}_{t},P^{T},t)_{j}$
23:end if
24:
b) Ablation Study of SPAA
For the ablation study, we randomly select 59 subjects including Pikachu, bed, cat, duck, handbag, table, microwave, pig, smartphone, Simba, bicycle, chair, elephant, hedgehog, monkey, pitcher, sofa, baseball, bird, chicken, fish, helmet, mouse, plate, spinach, basket, book, coat, football, horse, oven, rabbit, television, basketball, bread, cow, giraffe, lamp, panda, raccoon, tiger, bathtub, butterfly, cup, goat, laptop, pen, refrigerator, truck, bear, car, dog, hamster, lion, phone, sheep, van. We generate the image with the source template ”a photo of {}”, and the target template ”a photo of {}”. The attributes include both 43 types of colors and 14 types of materials. For different methods, we use the same random seed to generate the target attribute images. All target images undergo a two-step filtering process: LLaVa is used to eliminate incorrect attribute modifications, and the Dino Score is applied to remove images with substantial structural changes.
Figure 10: The Color Distribution of all subject in the Attribute-Dataset.
Figure 11: The Material Distribution of all subject in the Attribute-Dataset.
c) Attribute Editing Instruction Templates
1) List of color editing instruction templates for source to target image, where the first placeholder is the name of the subject, while the second placeholder is the name of the target color.
1) Change the color of the  {} to  {},
2) Change the color of  {} to  {},
3) Make the  {} appear  {},
4) Make the  {} to  {},
5) Turn the  {} to  {},
6) Modify the  {} to have a  {} appearance,
7) Make the  {} look  {},
8) Shift the base color of the  {} to  {},
9) Reconfigure the color of the  {} to achieve a  {},
10) Apply a color transformation to the  {}, making it  {},
11) Blend the  {} with a  {} color,
12) Recolor the  {} to  {},
13) Transform the  {} to a  {} color,
14) Enhance the  {} with a  {} color,
15) Adjust the  {} to reflect a  {} color,
16) Apply a color shift to the  {}, turning it  {},
17) Tone the  {} to a smooth  {} color,
2) List of color editing instruction templates for different color object target images, where the placeholder is the subject’s name, while the second and third placeholders are the original color name and target color name.
1) Change the color of the  {}  {} to  {},
2) Make the  {}  {} appear  {},
3) Make the  {}  {} to  {},
4) Turn the  {}  {} to  {},
5) Transform the  {}  {} to a  {} color,
6) Modify the  {}  {} to have a  {} appearance,
7) Recolor the  {}  {} to  {},
8) Make the  {}  {} look  {},
9) Alter the  {}  {} to a  {} shade,
10) Shift the hue of the  {}  {} to  {},
11) Enhance the  {}  {} with a  {} color,
3) List of material editing instruction templates, where the first placeholder is the name of the subject, while the second placeholder is the name of the target material.
1) Change the material of the {} to {},
2) Change the material of  {} to  {},
3) Transform the  {} to a  {} material,
4) Enhance the  {} with a  {} material,
5) Adjust the  {} to reflect a  {} material,
6) Apply a material shift to the  {}, turning it  {},
7) Apply a material transformation to the  {}, making it  {},
8) Retouch the  {} to have a  {} material scheme,
9) Refine the material of the  {} to a subtle  {},
10) Shift the base material of the  {} to  {},
11) Reconfigure the material of the  {} to achieve a  {},
4) List of material editing instruction templates for different material object target images, where the placeholder is the subject’s name, while the second and third placeholders are the original material name and target material name.
1) Change the material of the {} {} to {},
2) Transform the {} {} to a {} material,
3) Enhance the {} {} with a {} material,
4) Adjust the {} {} to reflect a {} material,
5) Apply a material shift to the {} {}, turning it {},
6) Apply a material transformation to the {} {}, making it {},
7) Retouch the {} {} to have a {} material scheme,
8) Refine the material of the {} {} to a subtle {},
9) Shift the base material of the {} {} to {},
10) Reconfigure the material of the {} {} to achieve a {},
d) Image Generation Prompt Templates, Attributes, Target Distribution
1) List of 19 prompt templates used to generate attribute-dataset:
1) a {}
2) a photo of {}
3) a picture of {}
4) a painting of {}
5) an image of {}
6) a toy of a {}
7) a rendition of a {}
8) a photo of a small {}
9) a close-up photo of a {}
10) a photo of a clean {}
11) a photo of a dirty {}
12) a cropped photo of the {}
13) a photo of the weird {}
14) a photo of a nice {}
15) a photo of the large {}
16) a photo of the small {}
17) a bright photo of the {}
18) a dark photo of the {}
19) a {} in the park
2) List of 43 color attributes for attribute-dataset:
amethyst, azure, beige, black, blue, bronze, brown, camel, copper, coral, cream, crimson, cyan, emerald, gold, gray, green, indigo, khaki, lime, magenta, maroon, navy, olive, orange, peach, pink, plum, purple, red, rose, salmon, silver, slate, tan, taupe, teal, tomato, turquoise, violet, white, wine, yellow
3) RGB Value of each color
amethyst: (153, 102, 204)
azure: (240, 255, 255)
beige: (245, 245, 220)
black: (0, 0, 0)
blue: (0, 0, 255)
bronze: (205, 127, 50)
brown: (165, 42, 42)
camel: (193, 154, 107)
copper: (184, 115, 51)
coral: (255, 127, 80)
cream: (255, 253, 208)
crimson: (220, 20, 60)
cyan: (0, 255, 255)
emerald: (80, 200, 120)
gold: (255, 215, 0)
gray: (128, 128, 128)
green: (0, 128, 0)
indigo: (75, 0, 130)
khaki: (195, 176, 145)
lime: (50, 205, 50)
magenta: (255, 0, 255)
maroon: (128, 0, 0)
navy: (0, 0, 128)
olive: (128, 128, 0)
orange: (255, 165, 0)
peach: (255, 218, 185)
pink: (255, 192, 203)
plum: (142, 69, 133)
purple: (128, 0, 128)
red: (255, 0, 0)
rose: (255, 0, 127)
salmon: (250, 128, 114)
silver: (192, 192, 192)
slate: (106, 128, 144)
tan: (210, 180, 140)
taupe: (72, 60, 50)
teal: (0, 128, 128)
tomato: (255, 99, 71)
turquoise: (64, 224, 208)
violet: (238, 130, 238)
white: (255, 255, 255)
wine: (159, 18, 54)
yellow: (255, 255, 0)
4) Target color and material attributes images number distribution: The distribution of the color editing dataset is given in Fig. 10 and the distribution of the material editing dataset is given in Fig. 11.
Figure 12: Results of Color alteration. As demonstrated, our method enables more precise color adjustments while effectively preserving the structural integrity of the subject.
Figure 13: Comparative results of color alteration with instruction-based editing methods on the most common seen colors. From top to bottom, the lamp color is changed to red, yellow, blue, green, and purple. Our method is capable of modifying the subject’s color while preserving its structural integrity and background information.
Figure 14: Comparative results of color alteration with instruction-based editing methods on less common seen colors. From top to bottom, the lamp color is changed to coral, plum, orange, tomato, and olive.
Figure 15: Comparative results of material alteration with instruction-based editing methods. From top to bottom, the dress material is changed to denim, leather, metal, plastic, and suede. Our method achieves the best balance between modifying the material and preserving the subject’s original structural integrity.
e) Qualitative Evaluation
For the color modification task, the most frequently observed colors such as red, yellow, blue, green, and purple, approaches like IP2P and HIVE may inadvertently modify the background while attempting to change the subject’s color (e.g., the table’s color is affected when altering the lamp to red, blue, or purple). Conversely, methods such as InsDiff, MGIE, and MagicBrush risk altering the subject’s structural integrity (e.g., when transforming the lamp’s color to red), see Fig. 13. For less frequently encountered colors, such as plum or tomato, methods like IP2P and HIVE may either fail to effectively change the subject’s color or unintentionally alter its structure, see Fig. 14. For the material changing task, our method strikes the optimal balance between material alteration and the retention of the original structure, see Fig. 15.
Figure 16: Screenshot from the human preference study. The user are asked to access which method is better terms of attribute modification, object structure preservation, background preservation, and overall editing quality.
f) Human Preference Study
As described in the main paper, we generate the images for the human preference study with the results of different instruction-based methods. To prevent potential bias, we use ”Method X” as a placeholder for the specific method. Object attribute editing refers to how well the object’s attributes are modified to the target attribute. Object structure preservation measures the extent to which the object’s original structure is retained in the edited image. The background preservation indicates how well the background remains intact without distortion. Overall editing quality encompasses the naturalness of the attribute modification, including the overall realism of the image post-editing. The format of the user study is provided in Fig. 16 where the user has to select between the results of six methods.

### v2

@@ 4 Object-Attribute Data Engine
Object-Attribute Editing instruction template. To construct the ATTRIBUTE dataset, initially, we collected class names from established datasets COCO [37], ImageNet [11], Open Images Dataset [33] and VisualGenome [32], resulting in a comprehensive list of more than 3,300 subjects, including foods, animals, plants, toys, clothing, transportations, daily necessities, natural objects , and more. Subsequently, for each subject, we generate the source images using 19 distinct prompt templates combined with randomly selected seeds. As altering a subject’s material inherently modifies its structure and texture, whereas color modification should minimally affect these properties, we divide the attribute dataset into a color-editing dataset and a material-editing dataset. To automatically generate the instructional text prompt for each pair of source-target images, we use the ability of GPT-4o [28] to generate a range of prompt template. Firstly, we randomly select 200 pairs of source and target images, where the object in the target images has varying attributes. Using the in-context learning ability of GPT-4o [28], we ask the model to identify the differences between each pair and to explain how we can convert one image to another image. Based on its responses, we develop two types of editing instruction templates to enhance attribute editing diversity: (a) instructions for transforming the source image into the target image, and (b) instructions for generating target images with different attributes. The first type of instruction guides the model in altering intrinsic object attributes to entirely different ones, while the second enables transitions between different attributes of the same object. For color modification, we add a new category of editing instruction templates focused on generating target images within the same color hue. This instruction enhances the model’s ability to fine-tune colors within the same hue by adjusting brightness and saturation levels. For each pair of source-target images, we randomly select one instruction template to create the triple training dataset. An example of a dataset sample is shown in Fig. 12.
5 InstructionAttribute
@@
Instructional Supervised Training. With the object attribute dataset, we adopt the instructional fine-tuning approach of InstructPix2Pix [3] to train our method to edit the attributes of the object. This approach utilizes a pre-trained Latent Diffusion Model (LDM) [53] as its primary framework, implementing a two-stage synthesis process that performs diffusion and denoising in a compressed latent space while ensuring high-quality results. During training, a pre-trained autoencoder with an encoder $E$ and a decoder $D$ is used to first map the edited image $x$ to its latent representation $z=E(x)$. The model then learns a network $\theta$ to predict the noise added to the noisy latent $z_{t}$, conditioned on both the image $c_{I}$ and the text instruction $c_{T}$, where $c_{T}$ is concatenated with $z_{t}$ and $\epsilon(c_{I})$. The network $\theta$ , which is initialized with Stable Diffusion weights, is fine-tuned to minimize the latent diffusion objective.
@@
Training Dataset. For color editing, we focus on 43 colors that are frequently observed in daily life, which including amethyst, azure, beige, black, blue, bronze, brown, camel, copper, coral, cream, crimson, cyan, emerald, gold, gray, green, indigo, khaki, lime, magenta, maroon, navy, olive, orange, peach, pink, plum, purple, red, rose, salmon, silver, slate, tan, taupe, teal, tomato, turquoise, violet, white, wine, yellow. In practice, we find that initially the cross-attention value amplifying ratio of the color descriptor to 5 and gradually reducing it by 0.1 per denoising step until it reaches 1.0 yields the optimal target color image success rate. Finally, we generate more than 0.82M pairs of source-target color image pairs after the image selection process. Through the instructional training dataset generation process, the color editing scenario, a triple dataset consisting of 1.8M input-output image pairs is generated, along with the corresponding edit instructions. For material editing, we examine 14 commonly observed materials in daily life, including cotton, glass, marble, plastic, velvet, denim, lace, mesh, wood, fur, leather, metal, suede, and wool. For the target material subject generated, we find that initializing the cross-attention Value amplifying ratio at 10 of the material descriptor and progressively reducing it by 0.2 per denoising step until reaching 1.0 achieves the highest success rate for generating target material images. In total, we generated over 0.31M source-target material image pairs after the image selection process. Using those image pairs, we create a triple dataset for the material editing scenario, consisting of 0.7M input-output image pairs and their corresponding editing instructions.
@@ 6 Experiment
Figure 13: Comparative Results of Instruction-Based Color Alteration on Common Colors. This figure presents a comparative analysis of color alteration capabilities using instruction-based editing methods. For the color modification task, InstructDiffusion (InsDiff) [15], MGIE [14], and MagicBrush [64] distorted the structure of the lamp. InstructPix2Pix (IP2P) [3] and HIVE [20] exhibited attribute leakage into the background area. HQ-Edit [27] altered the background and even removed the photo album. UltraEdit [67] failed to change the lamp’s color when attempting to modify it to red or blue. Our method effectively modifies the subject’s color while preserving both its structural integrity and background information.
Figure 14: Comparative Results of Instruction-Based Color Alteration in Less Common Hues. This figure presents a comparative analysis of color alteration capabilities using instruction-based editing methods on a selection of less commonly encountered colors. For the less common colors seen, InstructDiffusion (InsDiff) [15], MGIE [14], and MagicBrush [64] distorted the structure of the lamp. InstructPix2Pix (IP2P) [3] exhibited attribute leakage into the background area when changing the color to navy and turquoise. HQ-Edit [27] altered the background. HIVE [20] and UltraEdit [67] failed to change the lamp color when attempting to modify it to plum, navy, olive, turquois and teal.
Figure 15: Comparative Material Alteration with Instruction-Based Methods. This figure presents a comparative analysis of material alteration capabilities using various instruction-based editing methods. As can be seen, InstructPix2Pix (IP2P) [3], Instructdiffusion (InsDiff) [15], HIVE [20], MGIE [14], MagicBrush [64], HQ-Edit [27] and UltraEdit [67] distorted the structure of the dress. Instructdiffusion (InsDiff) [15], HIVE [20], UltraEdit [67] exhibited attribute leakage into the background. Our method achieves the optimal balance between modifying the material and preserving the subject’s original structural.
Figure 16: Results of attribute alterations using our InstructAttribute. a) This figure demonstrates the ability of our method in performing precise color adjustments. b) This figure demonstrates our method’s capability to modify an object’s color while strictly maintaining its hue. c) This figure illustrates the results of material modification using our method.
@@
Baseline models. We make comparisons with the latest instruction-based image editing methods, including InstructPix2Pix (IP2P) [3], Instructdiffusion (InsDiff) [15], HIVE [20], MGIE [14], MagicBrush [64], HQ-Edit [27] and UltraEdit [67]. To maintain reproducibility and fairness, we employed the default hyper-parameters provided in the official implementations.
Qualitative evaluation. For the color modification task, as illustrated in Fig.13, prevalent colors such as red, yellow, blue, green, and purple, methods such as IP2P and HIVE may unintentionally alter the color of the background elements while adjusting the subject’s color (e.g., the color of the table is affected when changing the lamp to red, blue, or purple), while HQ-Edit change the structure of background area, for example, it remove the books and album when change the color of the lamp to red and yellow. Methods such as InsDiff, MGIE, and MagicBrush may compromise structural coherence during color transformations (e.g., when the lamp’s color is transformed to red), leading to unintended geometric distortions. Although UltraEdit demonstrates superior structural preservation, its ability to achieve precise color transitions for target objects remains inconsistent, frequently failing to meet desired chromatic outcomes. In contrast, our InstructAttribute achieves superior precision in color modification while preserving both the structural integrity of objects and the consistency of background elements, ensuring that no unintended alterations occur during the editing process. Our method achieves consistent performance even for less common colors, such as plum, navy, olive, turquoise, and teal, as demonstrated in Fig. 14. For material modification tasks, as demonstrated in Fig. 15, current instructed-based methodologies, including IP2P, InsDiff, HIVE, MGIE, MagicBrush, HQ-Edit, and UltraEdit, exhibit limitations in the execution of precise material transformations while frequently inducing unintended structural distortions or attribute leakage to the background area. In contrast, our proposed approach achieves a refined equilibrium between the modification of the target material and the rigorous preservation of the geometric and structural fidelity of the object, establishing itself as the benchmark for precision in comparative evaluations. Furthermore, Fig. 16 a) shows that our model is capable of accurate and precise color adjustments while maintaining the structural integrity of the subject. Fig. 16 b) reveals that our method is able to adjust an object’s color within the same hue by modifying its saturation and brightness, preserving the original hue while achieving targeted chromatic variations. Fig. 16 c) illustrates that our method ensures a consistent effect of material transformation on various subjects while preserving the structure of the object and the background information.
@@
IP2P 14.2% 7.4% 6.2% 7.1%
InsDiff 4.1% 3.9% 4.5% 1.8%
HIVE 8.1% 7.6% 6.3% 6.5%
MGIE 5.8% 2.1% 3.2% 3.8%
MagicBrush 10.2% 4.3% 7.6% 7.2%
HQ-Edit 7.9% 2.3% 1.2% 4.3%
UltraEdit 3.3% 5.6% 7.6% 6.4%
Ours 46.4% 66.8% 63.4% 62.9%
@@
Quantitative evaluation. Given the lack of publicly available datasets for evaluating material and color modification tasks, we developed two specialized datasets designed for quantitative analysis, addressing material and color modification tasks independently. For the material modification task, we randomly selected 100 subjects and generated 239 source images with StableDiffusion [53] checkpoints V1.4 with a random seed. For each source image, we modified the material to all 14 fine-grained materials, resulting in a total of 3,346 pairs of source and target images. For evaluation metrics, we used the DINO score [6] on a gray scale to measure semantic similarity, SSIM [62] to measure structural similarity, and used the CLIP Score (CS) [51] to measure similarity between the target image and its corresponding text description. For the color modification task, since any object can undergo a color change, we enlarge the subject type to 659 and generate a source image for each subject using the StableDiffusion [53] checkpoints V1.4. For each source image, we used the SAM method [31] to generate the object mask and manually selected the best, resulting in 659 source images with the corresponding subject mask. For each source image, we modified the color to all 43 fine-grained colors, resulting in a total of 28,337 pairs of source and target images. For evaluation metrics, except for using the DINO score, SSIM, and CLIP score, we used LPIPS [65] in the background area to measure the preservation of background information. L1 loss of HVS of the object area between the target image and the corresponding pure color image to measure how well the color of the object has been changed. As can be seen in Table 2, for color alteration, our method performs best in preserving structure, background information, and the best in changing the object’s color compared to other instruct-based methods. For the material modification task, our method performs best in preserving the structure while having the highest text-image similarity after we changed the material of the subject.
@@
Human Preference Study. We conduct a human preference study to measure the performance of different methods asking: (1) which method produces the best attributes modification; (2) which method produces the best preserves the structure of the subject; (3) which method produces the best preserves the background; (4) which method produces the best overall attribute editing quality. We randomly sampled 100 source-target examples and let 20 annotators assess those images. Object attribute editing refers to how well the object’s attributes are modified to the target attribute. Object structure preservation measures the extent to which the object’s original structure is retained in the edited image. The background preservation indicates how well the background remains intact without distortion. Overall editing quality encompasses the naturalness of the attribute modification, including the overall realism of the image post-editing. The results of the human preference study are listed in Table 3. As can be seen, our method performs the best in all aspects, and the results are consistent with the quantitative experiments.
@@ 7 Conclusion
We propose an instruction-based model for object-level color and material modification, enabling precise control over visual attributes while preserving structural and contextual integrity. First, we introduce a training-free Structure Preservation and Attribute Amplification (SPAA) method to generate diverse color/material variations without additional fine-tuning. Using SPAA, we curate the Attribute Dataset, a large-scale resource for object editing, encompassing varied objects, colors, and materials. Training our model on this dataset achieves an optimal balance between accurate attribute editing and preservation of object structure and background. Limitations and Future Work. While SPAA produces plausible variations, cross-attention leakage occasionally causes misaligned edits. Current filtering via LLaVA, Dino, Grounding Dino, and SAM risks excluding a valid target image. Our future work will refine attention mechanisms to constrain focus to target objects, mitigating interference from irrelevant regions.